OpenStudy (aem9296):

4. A jet engine is moving an airplane forward at a speed of 900 km/hr. If the thrust of the engine is 1 x 105 N, what is the power developed by the engine? a. 5 MW b. 10 MW c. 25 MW d. 50 MW

OpenStudy (deoxna):

Power is a measure of energy input/output per unit time, or Joules per second. Mechanical power can also relate force (thrust) and velocity by: \[ P=F*v\] The only problem is that you velocity is given in km/h, which are not standard units (Watts are in J/s). So convert km/h to m/s: \[900 km/h \div 3.6 = 250m/s\] \[250m/s*10^5N=25*10⁶W\] or 25 megawatts.

OpenStudy (aem9296):

Thanks so much!!!

OpenStudy (deoxna):

My pleasure!